A fair die is rolled 10 times. What is the average number of even number outcomes?
Answer #1
Concepts and reason

Binomial distribution:

In a series of n independent trials, for each trial if the probability of success is a constant p and the probability of failure is q=1-p
; then, the probability of x success and obviously (x-u)
failures is given by the Binomial distribution.

Requirements of a binomial experiment are as follows:

• For the given experiment, each trial has two possibilities. That is, it has a success and a failure, which are mutually exclusive outcomes.

• The number of trials in the given experiment is known in advance, and they are fixed.

• The outcomes of the trials are independent.

• The probability of success is unaltered for each trail throughout the experiment.

Fundamentals

Binomial probability of success:

p=
number of success
total number of trials

Where n is the number of trials and p is the probability of success for each trial.

Binomial probability of failure:

p=1-9

Where,

9: probability of failure
.

Binomial distribution probability mass function:

The binomial distribution is the discrete probability distribution and its probability is,

P(X = x) =
p)*-* ;Here x = 0,1,2,...,n for 0 spsi

Also, x! n-x)!

Where n is the number of trials and p is the probability of success for each trial.

• The mean formula for the binomial distribution is, du = 11
.

• The standard deviation formula for the binomial distribution is, o = Jnp(1- p)
.

The objective of the problem is determined below:

From the information given, a fair die is rolled 10 times. That is, n=10
, and the probability of the outcome is an even number is 0.5. That is, p=0.5
, by using the probability and sample size the average number of even number outcomes is obtained.

The average number of even number outcomes is obtained as shown below:

The expected value is,

u=np
= 10x0.50
= 5

Ans:

The average number of even number outcomes is 5.